17. Evaluate the integral. 3 y3 - 2y2 - y dy y2
Answer #1

\text{Let } I=\int_1^3\frac{y^3-2y^2-y}{y^2}dy

=> I=\int_1^3\left (\frac{y^3}{y^2} -\frac{2y^2}{y^2}-\frac{y}{y^2} \right )dy =\int_1^3\left (y -2-\frac{1}{y} \right )dy

=> I=\int_1^3ydy -2\int_1^3dy-\int_1^3\frac{1}{y}dy

\left ( \because \int_{a}^{b} (k_1f(x)+ k_2 g(x))dx = k_1\int_{a}^{b}f(x)dx + k_2 \int_{a}^{b} g(x)dx\;\:\text{Where } k_1 , k_2 \text{ are constants. }\right )

=> I=\left ( \frac{y^2}{2} \right )_{y \:=\:1}^{y\:=\:3} -2\left (y \right )_{y \:=\:1}^{y\:=\:3}-\left (ln\left | y \right | \right )_{y\:=\:1}^{y\:=\:3}

=> I=\left ( \frac{(3)^2}{2} -\frac{(1)^2}{2} \right ) -2\left (3-1 \right )-\left (ln\left | 3 \right | -ln\left | 1 \right | \right )

=> I=\left ( \frac{9}{2} -\frac{1}{2} \right ) -4-\left (ln\left | 3 \right | -0 \right )

=> I=\frac{8}{2} -4-ln\left | 3 \right |

=> I=4 -4-ln\left | 3 \right |

=> I=-ln\left | 3 \right | =-ln(3)

=> I=- 1.098612